# ConspEmoLLM: Conspiracy Theory Detection Using an Emotion-Based Large Language Model

### v1

@@ 3.2.4 Construction of the ConDID conspiracy detection instruction dataset
Task
Raw (Train/Dev/Test)
Instruction (Train/Dev/Test)
Task 1
2092/697/698
Task 2
2092/697/698
2092/697/698
Task 3
25104/8364/8376
Task 4
669/223/233
Task 5
669/223/233
669/223/233
@@
Task
Prompt Template
Task 1
Task: Classify the text regarding COVID-19 conspiracy theories or misinformation into one of the following three classes: 0. Unrelated. 1. Related (but not supporting). 2. Conspiracy (related and supporting).
Task 2
Task: detect whether the text in any form mentions or refers to any of the specific categories of COVID-19 conspiracy theories (’suppressed cures’, ’behavior control’, ’anti vaccination’, ’fake virus’, ’intentional pandemic’, ’harmful radiation’, ’depopulation’, ’new world order’, ’satanism’, ’esoteric misinformation’, ’other conspiracy theory’) or other misinformation. If it doesn’t, it is ’no conspiracy.
Task 3
Task: Classify the text regarding the specific category [Specific Conspiracy] into one of the following three classes: 0. Unrelated. 1. Related (but not supporting). 2. Conspiracy (related and supporting).
Task 4
Task: Determine if the text is a conspiracy theory. Classify it into one of the following two classes: 0. non-conspiracy. 1. conspiracy.
Task 5
Task: Determine the relatedness between the text and [conspiracy theory]. Classify it into one of the following three classes: 0. not related. 1. closely related. 2. broadly related.
Affective prompt
Task: original task prompt + "You can also refer to the affective information. (1) Emotion intensity: anger: 0.521, fear: 0.625, joy: 0.25, sadness: 0.354. (2) Ordinal classification of emotion intensity: moderate amount of anger can be inferred. low amount of fear can be inferred. no joy can be inferred. no sadness can be inferred. (3) Sentiment intensity: 0.435. (4) Sentiment classification: neutral or mixed mental state can be inferred. (5) The emotions included are: anger, disgust, fear."
@@
We used the raw dataset as the basis to build the instruction dataset. We randomly selected 20% of the data as the test set and 20% as the validation set. The dataset statistics are presented in Table 1. For Task 3, prompts needed to be created for each of the 12 categories, resulting in a magnitude of data that is 12 times the size of the original corpus. We constructed instruction-tuning data for each task based on the following template:
Task: [task prompt] Text: [Text] Class/The text mentions or refers to conspiracies: [output]
[task prompt] denotes the instruction for the task. [input text] is a data item from the raw data. With the exception of Task 2, all tasks use Class to prompt LLM to generate answers. [output] is the output from LLM. Table 2 lists the task prompts for each task, and Figure 6 presents several examples used to fine-tune the LLM. To allow an evaluation of the impact of explicitly encouraging the LLMs to make use of affective information, we additionally constructed explicit prompts of the form shown in the final row of Table 2. These prompts provide information about the analysis results obtained from EmoLLaMA regarding the text to be classified.
@@ 3.3 ConspEmoLLM and ConspLLM
We built ConspEmoLLM by fine-tuning EmoLLaMA-chat-7b Liu et al. (2024a) using the ConDID dataset. We also fine-tuned an LLM that does not use affective information (ConspLLM) based on LLaMA2-chat-7b Touvron et al. (2023), also using ConDID. The models are trained based on the AdamW optimizer Loshchilov and Hutter (2017) for three epochs, using DeepSpeed Rasley et al. (2020) to reduce memory usage. We set the batch size to 256. The initial learning rate is set to 1e-6 with a warm-up ratio of 5%, and the maximum model input length is set to 4096. All models are trained on two Nvidia Tesla A100 GPUs, each with 80GB of memory. Figure 6 provides an overview of multi-task instruction tuning of ConspEmoLLM for diverse conspiracy detection tasks.
4 Experiments
4.1 Baseline models
PLMs: Conspiracy theory detection is typically regarded as a classification task. For our baseline models, we selected commonly used PLMs, which can only be fine-tuned for individual tasks, i.e., the general language BERT and RoBERTa, along with CT-BERT Müller et al. (2023), which is tailored for the COVID-19 domain. We treat Tasks 1 and 5 as 3-way classification tasks, and Task 4 as a binary classification task, using cross-entropy loss for training. Task 2 is treated as a multi-label binary classification problem, which can be achieved by utilizing the binary cross-entropy with logits loss. To address Task 3, fine-tuning is performed for 12 different classification problems, each with its own cross-entropy loss function. The final loss is the average of the 12 losses.
LLMs: LLMs have been proven to be capable of solving numerous tasks. We apply zero-shot prompting on the instruction dataset to the following open-source LLMs: Falcon-7B-instruct Penedo et al. (2023), LLaMA2-chat-7B Touvron et al. (2023), OPT-7B Zhang et al. (2022), BLOOM-7B Workshop et al. (2022), and Vicuna-7B-v1.5333https://huggingface.co/lmsys/vicuna-13b-v1.5. We also utilize zero-shot prompting with the proprietary LLM ChatGPT.
4.2 Evaluation methods
4.3 Results
@@
ACC
PRE
REC
F1
ACC
PRE
REC
F1
ACC
PRE
REC
F1
BERT
0.576
0.601
0.428
0.406
0.272
0.150
0.012
0.023
0.893
0.871
0.893
0.842
RoBERTa
0.517
0.505
0.335
0.231
0.279
0.204
0.077
0.112
0.893
0.842
0.893
0.844
CT-BERT
0.564
0.472
0.469
0.428
0.042
0.137
0.303
0.175
0.893
0.797
0.893
0.842
Falcon
0.265
0.361
0.265
0.211
0.189
0.193
0.180
0.160
0.556
0.810
0.556
0.653
Vicuna
0.380
0.446
0.380
0.392
0.136
0.172
0.270
0.150
0.437
0.825
0.437
0.556
LLaMA2-chat
0.325
0.527
0.325
0.251
0.193
0.093
0.071
0.074
0.674
0.814
0.674
0.732
OPT
0.311
0.389
0.311
0.298
0.235
0.216
0.048
0.072
0.481
0.805
0.481
0.590
BLOOM
0.328
0.389
0.328
0.318
0.268
0.000
0.000
0.000
0.114
0.810
0.114
0.146
ChatGPT
0.638
0.677
0.638
0.596
0.324
0.546
0.333
0.332
0.208
0.896
0.208
0.240
ChatGPT-aff
0.583
0.647
0.583
0.525
0.312
0.449
0.326
0.308
0.150
0.898
0.150
0.146
ConspLLM
0.662
0.757
0.662
0.675
0.328
0.685
0.320
0.334
0.893
0.886
0.893
0.864
ConspLLM-aff
0.517
0.483
0.517
0.354
0.203
0.404
0.330
0.301
0.077
0.901
0.077
0.015
ConspEmoLLM 0.695
0.755
0.695 0.705 0.340 0.699 0.345 0.364 0.897
0.884
0.897
0.860
@@
Tables 3 and 4 report the results for each task. ChatGPT-aff and ConspLLM-aff are the models that use the prompts containing explicit affective information. From the results tables, we can observe that, in terms of F1 score, the fine-tuned ConspLLM and ConspEmoLLM outperform all other open-source models [^4] It should be noted that the LLMs that are not instruction-tuned produced some responses that do not follow instructions, i.e., they do not produce output of the type requested in the prompts. In such cases, we label them as unrelated or non-conspiracy, depending on the task., as well the PLMs. ConspLLM and ConspEmoLLM also both outperform ChatGPT, with the exception of ConspLLM on Task 4. Moreover, ConspEmoLLM, which is fine-tuned based on a large emotion language model, achieves F1 scores that are over 3% higher than ConspLLM for all tasks except for Task 3, in which it lags slightly behind ConspLLM. The results for ChatGPT-aff and ConspLLM-aff reveal that explicitly augmenting prompts with affective information leads to reduced performance for all Tasks apart from Task 4. We can infer that explicitly adding affective information appears to distract models’ attention from the task at hand. In contrast, the implicit use emotion information by ConspEmoLLM is more successful in allowing emotional cues to be leveraged.
@@ 5 Conclusion
In this paper, our comprehensive affective analysis of two conspiracy theory datasets demonstrated that conspiracy theory text exhibits sentiment and emotion features that are distinct from mainstream text. The results of this analysis motivated our development of ConspEmoLLM, an open-source domain-specific LLM that is based on affective information, and which can perform diverse conspiracy theory detection tasks. ConspEmoLLM was fine-tuned using our newly constructed multitask conspiracy detection instruction dataset (ConDID). Evaluation on the test set of ConDID reveals that ConspEmoLLM achieves SOTA performance among the other open-source LLMs tested, as well as ChatGPT in all tasks. Its performance on most tasks surpasses that of the ConspLLM model, which was also instruction-tuned using ConDID, but does not use affective information. These results provide strong evidence of the importance of affective features in detecting various types of information relating to conspiracy theories.
As future work, we will augment the ConDID dataset with further conspiracy theory datasets, including data from multiple platforms, sources, domains and languages. This should help to further improve the performance and diversity of tasks that can be carried out using ConspEmoLLM. We will additionally explore alternative methods of incorporating affective information to further improve the ability of the model to detect conspiracy theories. Furthermore, we will design more appropriate prompts and utilize more complex and diverse model structures to better leverage emotions and sentiments.
@@ Acknowledgements
Appendix A Appendix
A.1 Affective analysis for different conspiracies on COCO dataset (from Figure 7 to Figure 11)
Figure 7: Emotion intensity of different conspiracies
Figure 8: Ordinal classification of emotion intensity of different conspiracies
Figure 9: Sentiment strength of different conspiracies
Figure 10: Sentiment classification of different conspiracies
Figure 11: Emotion classification of different conspiracies
A.2 Affective analysis on LocoAnnotations (from Figure 12 to Figure 21)
Figure 12: Emotion intensity of conspiracy and non-conspiracy
Figure 13: Ordinal classification of emotion intensity of conspiracy and non-conspiracy
Figure 14: Sentiment strength of conspiracy and non-conspiracy
Figure 15: Sentiment classification of conspiracy and non-conspiracy
Figure 16: Emotion classification of conspiracy and non-conspiracy
Figure 17: Emotion intensity of different relatedness
Figure 18: Ordinal classification of emotion intensity of different relatedness
Figure 19: Sentiment strength of different relatedness
Figure 20: Sentiment classification of different relatedness
Figure 21: Emotion classification of different relatedness

### v3

@@ 3.2.4 Construction of the ConDID conspiracy detection instruction dataset
Task Raw (Train/Dev/Test) Instruction (Train/Dev/Test)
Task 1 2092/697/698
Task 2 2092/697/698 2092/697/698
Task 3 25104/8364/8376
Task 4 669/223/233
Task 5 669/223/233 669/223/233
@@
Task Prompt Template
Task 1 Task: Classify the text regarding COVID-19 conspiracy theories or misinformation into one of the following three classes: 0. Unrelated. 1. Related (but not supporting). 2. Conspiracy (related and supporting).
Task 2 Task: detect whether the text in any form mentions or refers to any of the specific categories of COVID-19 conspiracy theories (’suppressed cures’, ’behavior control’, ’anti vaccination’, ’fake virus’, ’intentional pandemic’, ’harmful radiation’, ’depopulation’, ’new world order’, ’satanism’, ’esoteric misinformation’, ’other conspiracy theory’) or other misinformation. If it doesn’t, it is ’no conspiracy.
Task 3 Task: Classify the text regarding the specific category [Specific Conspiracy] into one of the following three classes: 0. Unrelated. 1. Related (but not supporting). 2. Conspiracy (related and supporting).
Task 4 Task: Determine if the text is a conspiracy theory. Classify it into one of the following two classes: 0. non-conspiracy. 1. conspiracy.
Task 5 Task: Determine the relatedness between the text and [conspiracy theory]. Classify it into one of the following three classes: 0. not related. 1. closely related. 2. broadly related.
Affective prompt Task: original task prompt + "You can also refer to the affective information. (1) Emotion intensity: anger: 0.521, fear: 0.625, joy: 0.25, sadness: 0.354. (2) Ordinal classification of emotion intensity: moderate amount of anger can be inferred. low amount of fear can be inferred. no joy can be inferred. no sadness can be inferred. (3) Sentiment intensity: 0.435. (4) Sentiment classification: neutral or mixed mental state can be inferred. (5) The emotions included are: anger, disgust, fear."
@@
We used the raw datasets as the basis to build the instruction dataset. We randomly selected 20% of the data as the test set and 20% as the validation set. The dataset statistics are presented in Table 1. For Task 3, prompts needed to be created for each of the 12 categories, resulting in a magnitude of data that is 12 times the size of the original corpus. We constructed instruction-tuning data for each task based on the following template:
Task: [task prompt] Text: [input text] Class/The text mentions or refers to conspiracies: [output]
[task prompt] denotes the instruction for the task. [input text] is a data item from the raw data. With the exception of Task 2, all tasks use Class to prompt LLM to generate answers. [output] is the output from LLM. Table 2 lists the task prompts for each task, and Figure 6 presents several examples used to fine-tune the LLM. To allow an evaluation of the impact of explicitly encouraging the LLMs to make use of affective information, we additionally constructed prompts of the form shown in the final row of Table 2. These prompts provide information about the specific analysis results obtained from EmoLLaMA regarding the text to be classified.
@@ 3.3 ConspEmoLLM and ConspLLM
We built ConspEmoLLM by fine-tuning EmoLLaMA-chat-7b [13] using the ConDID dataset. We also fine-tuned an LLM that does not use affective information (ConspLLM) based on LLaMA2-chat-7b [27], also using ConDID. The models are trained based on the AdamW optimizer [15] for three epochs, using DeepSpeed [25] to reduce memory usage. We set the batch size to 256. The initial learning rate is set to 1e-6 with a warm-up ratio of 5%, and the maximum model input length is set to 4096. All models are trained on two Nvidia Tesla A100 GPUs, each with 80GB of memory. Figure 6 provides an overview of multi-task instruction tuning of ConspEmoLLM for diverse conspiracy detection tasks.
@@
ACC PRE REC F1 ACC PRE REC F1 ACC PRE REC F1
BERT 0.576 0.601 0.428 0.406 0.272 0.150 0.012 0.023 0.893 0.871 0.893 0.842
RoBERTa 0.517 0.505 0.335 0.231 0.279 0.204 0.077 0.112 0.893 0.842 0.893 0.844
CT-BERT 0.564 0.472 0.469 0.428 0.042 0.137 0.303 0.175 0.893 0.797 0.893 0.842
Falcon 0.265 0.361 0.265 0.211 0.189 0.193 0.180 0.160 0.556 0.810 0.556 0.653
Vicuna 0.380 0.446 0.380 0.392 0.136 0.172 0.270 0.150 0.437 0.825 0.437 0.556
LLaMA2-chat 0.325 0.527 0.325 0.251 0.193 0.093 0.071 0.074 0.674 0.814 0.674 0.732
OPT 0.311 0.389 0.311 0.298 0.235 0.216 0.048 0.072 0.481 0.805 0.481 0.590
BLOOM 0.328 0.389 0.328 0.318 0.268 0.000 0.000 0.000 0.114 0.810 0.114 0.146
ChatGPT 0.638 0.677 0.638 0.596 0.324 0.546 0.333 0.332 0.208 0.896 0.208 0.240
ChatGPT-aff 0.583 0.647 0.583 0.525 0.312 0.449 0.326 0.308 0.150 0.898 0.150 0.146
ConspLLM 0.662 0.757 0.662 0.675 0.328 0.685 0.320 0.334 0.893 0.886 0.893 0.864
ConspLLM-aff 0.517 0.483 0.517 0.354 0.203 0.404 0.330 0.301 0.077 0.901 0.077 0.015
ConspEmoLLM 0.695 0.755 0.695 0.705 0.340 0.699 0.345 0.364 0.897 0.884 0.897 0.860
@@
4 Experiments
4.1 Baseline models
PLMs: Conspiracy theory detection is typically regarded as a classification task. For our baseline models, we selected commonly used PLMs, which can only be fine-tuned for individual tasks, i.e., the general language BERT and RoBERTa, along with CT-BERT [18], which is tailored for the COVID-19 domain. We treat Tasks 1 and 5 as 3-way classification tasks, and Task 4 as a binary classification task, using cross-entropy loss for training. Task 2 is treated as a multi-label binary classification problem, which can be achieved by utilizing the binary cross-entropy with logits loss. To address Task 3, fine-tuning is performed for 12 different classification problems, each with its own cross-entropy loss function. The final loss is the average of the 12 losses.
LLMs: LLMs have been proven to be capable of solving numerous tasks. We apply zero-shot prompting on the instruction dataset to the following open-source LLMs: Falcon-7B-instruct [21], LLaMA2-chat-7B [27], OPT-7B [35], BLOOM-7B [28], and Vicuna-7B-v1.5333https://huggingface.co/lmsys/vicuna-13b-v1.5. We also utilize zero-shot prompting with the proprietary LLM ChatGPT.
4.2 Evaluation methods
4.3 Results
Tables 3 and 4 report the results for each task. ChatGPT-aff and ConspLLM-aff are the models that use the prompts containing explicit affective information. From the results tables, we can observe that, in terms of F1 score, the fine-tuned ConspLLM and ConspEmoLLM outperform all other open-source models [^4] It should be noted that the LLMs that are not instruction-tuned produced some responses that do not follow instructions, i.e., they do not produce output of the type requested in the prompts. In such cases, we label them as unrelated or non-conspiracy, depending on the task., as well the PLMs. ConspLLM and ConspEmoLLM also both outperform ChatGPT, with the exception of ConspLLM on Task 4. Moreover, ConspEmoLLM, which is fine-tuned based on a large emotion language model, achieves F1 scores that are over 3% higher than ConspLLM for all tasks except for Task 3, in which it lags slightly behind ConspLLM. The results for ChatGPT-aff and ConspLLM-aff reveal that explicitly augmenting prompts with affective information leads to reduced performance for all tasks apart from Task 4. We can infer that explicitly adding affective information appears to distract models’ attention from the task at hand. In contrast, the implicit use of emotion information by ConspEmoLLM is more successful in allowing emotional cues to be leveraged.
5 Automatic predictions
The following example code demonstrates how ConspEmoLLM can be applied to automatically classify text according to whether or not it represents a conspiracy theory. Any piece of text may be classified by ConspEmoLLM, by replacing [input text] with the text of interest. The task prompt may also be adjusted to perform the other conspiracy-related tasks introduced above. Further details can be found on Github  [^5] https://github.com/lzw108/ConspEmoLLM.
1from transformers import AutoTokenizer, AutoModelForCausalLM
2Model_PATH = "lzw1008/ConspEmoLLM-7b"
3tokenizer = AutoTokenizer.from_pretrained(MODEL_PATH)
4model = AutoModelForCausalLM.from_pretrained(MODEL_PATH, device_map=’auto’)
5
6prompt = ’’’Human:
7Task:Task:Determineifthetextisaconspiracytheory.Classifyitintooneofthefollowingtwoclasses:0.non-conspiracy.1.conspiracy
8Text:[inputtext]
9Class:
10Assistant:
11’’
12inputs = tokenizer(prompt, return_tensors="pt")
13generate_ids = model.generate(inputs["input_ids"], max_length=256)
14response = tokenizer.batch_decode(generate_ids, skip_special_tokens=True)[0]
15print(response)
16>> 0. non-conspiracy / 1. conspiracy
@@ 6 Conclusion
In this paper, our comprehensive affective analysis of two conspiracy theory datasets has demonstrated that conspiracy theory text exhibits sentiment and emotion features that are distinct from mainstream text. The results of this analysis motivated our development of ConspEmoLLM, an open-source domain-specific LLM that is based on affective information, and which can perform diverse conspiracy theory detection tasks. ConspEmoLLM was fine-tuned using our newly constructed multitask conspiracy detection instruction dataset (ConDID). Evaluation on the test set of ConDID reveals that ConspEmoLLM achieves SOTA performance among the other open-source LLMs tested, as well as ChatGPT, in all tasks. On most tasks, the performance of ConspEmoLLM surpasses that of the ConspLLM model, which was also instruction-tuned using ConDID, but does not use affective information. These results provide strong evidence of the importance of affective features in detecting various types of information relating to conspiracy theories. We also exemplified how ConspEmoLLM can be applied straightforwardly to detect the presence of conspiracy-related information in any piece of text. This demonstrates the real world utility of our model in helping people to identify the misinformation on the Internet, thus contributing towards reducing the potential harm caused by conspiracy theories.
As future work, we aim to augment the ConDID dataset with further conspiracy theory datasets, including data from multiple platforms, sources, domains and languages. This should help to further improve the performance and diversity of tasks that can be carried out using ConspEmoLLM. We will additionally explore alternative methods of incorporating affective information to further improve the ability of the model to detect conspiracy theories. Furthermore, we will design more appropriate prompts and utilize more complex and diverse model structures to better leverage emotions and sentiments.